Identify the greater number, wherever possible, in each of the following?
2^8 or 8^2

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The expansion of 2^8= 2 × 2 × 2 × 2 × 2 × 2 × 2 × 2 = 256
The expansion of 8^2= 8 × 8= 64
Clearly,
256 > 64
So, 2^8> 8^2
Hence 2^8is thegreater number.

Thanks